# Signage-Aware Exploration in Open World using Venue Maps

### v1

@@ IV-C Exploration-Exploitation Planning
Candidate Next Pose: During navigating to the next landmark $g_{k}$, the candidate next poses $c\in V$ are sampled among the set of unvisited frontiers $\mathcal{F}$ and unvisited viewpoints $\mathcal{V}$. The frontiers $f\in\mathcal{F}$ are sampled at the boundaries between known and unknown spaces, which biases the exploration to unknown regions. The viewpoints $v\in\mathcal{V}$ are sampled nearby the signage, allowing the robot to approach the signage and adjust its orientation for better recognition.
Next Best Pose Selection: We tailor the frontier utility $U_{f}$ and viewpoint utility $U_{v}$ for the decision on the candidate selection, respectively. At the initial stage, we induce the robot to explore its surroundings by selecting the frontiers that maximize the information gain $\sum G$ within the camera’s FOV. After at least two landmarks are found, we align the online map with venue map and estimate all the landmark poses $p_{g_{t}}$ at the world coordinate using random sample consensus (RANSAC) [35] and iterative closest points (ICP) [36]. Therefore, we perform the frontier-based informed search using the relative directions between landmarks on the venue maps. To this end, a directional heuristic is calculated to favor selecting the frontiers closest to the direction towards the subgoal $g_{t}$ (see Fig. 4), which is defined as the normalized inner product between the vectors from the current position $p_{t}$ to the frontier $f_{i}$ and the vector from the current position $p_{t}$ to the position of next subgoal $p_{g_{t}}$. Specifically, we first calculate the transformation matrix $T=[R,t]$ and scale $\alpha$, which is updated whenever a new landmark is found. Then, we back-project the landmark positions from the venue map to the world coordinate. As such, the directional heuristic $h(f_{i})$ is calculated as $h(f_{i})=\frac{(p_{t}-f_{i})\cdot(p_{t}-T^{-1}p_{g_{t}})}{\|p_{t}-f_{i}\|*\|p_{t}-T^{-1}p_{g_{t}}\|}$, and thereby the frontier utility $U_{f}(f_{i},p_{t})$ is computed as:
Figure 4: We estimate the position of next landmark by calculating coordinate transformation between online map and venue map for guiding exploration.
@@
where $s_{v}(v_{j})=s_{\phi}(o_{\tau_{j}},o_{\tau_{j}}^{\prime})$ denotes the score that the signage $\tau_{j}$ belonging to $v_{j}$ is matched to the landmarks on venue map, and $d(p_{t},v_{j})$ is the navigation cost with a weight $\eta$. A factor $\beta$ that determines to what extent of scores we should highlight the viewpoint candidates than frontiers is employed to handle the exploration-exploitation dilemma. Intuitively, the priority of exploiting signage originates from the fact that it can help localize the robot on the venue maps and determine the decision on the next best pose. Finally, we obtain the optimal next pose $c_{t+1}$ among unvisited frontiers and viewpoints that maximizes the overall utility $U(c_{j},p_{t})$:
@@
where $c_{j}\in\mathcal{F}\cup\mathcal{V}$ and $I_{f}(c_{j})=1\ \text{if}\ c_{j}\in\mathcal{F}\ \text{else}\ 0$. Then we plan a collision-free path to $c_{t+1}$ using the information RRT* planner from Open Motion Planning Library [37] with a fixed replan rate, and follow the route by a local planner from autonomous exploration development environment [^1] https://www.cmu-exploration.com/, which supports dynamic collision avoidance in the scenes. The overall procedure is summarized in algorithm 1.
@@
0:  Venue map $\mathbf{M}$ including landmark names $\mathcal{T}$ and a topological graph $\mathcal{G}$, pseudo text images $\{o^{\prime}\}$, text spotter $\phi$, RGB $o_{t}$, depth $d_{t}$, LiDAR $\mathcal{X}_{t}$
1:  $\mathcal{M}\leftarrow\emptyset$, $\mathcal{F}\leftarrow\emptyset$, $\mathcal{V}\leftarrow\emptyset$
2:  while not done do
3:     $o_{t},d_{t},\mathcal{X}_{t}\leftarrow$ getObservation()
4:     ${\mathcal{M}},f_{t}\leftarrow$ mappingModule($\mathcal{X}_{t}$)
5:     $\mathcal{F}\leftarrow\mathcal{F}\cup f_{t}$
6:     $R^{2D},f^{2D},v_{t}\leftarrow\phi(o_{t})$
7:     if $R^{2D}$ is not None then
8:        $R^{3D},f^{3D}\leftarrow$ Fusion($R^{2D},f^{2D},d_{t},\mathcal{M}$)
9:        $\tau_{t}\leftarrow$ textInstanceRetrieval($f^{3D},\{o^{\prime}\}$)
10:        $g_{k}\leftarrow$ topologicalPlanning($\tau_{t},\mathcal{G}$)
11:        $\mathcal{V}\leftarrow\mathcal{V}\cup v_{t}$
12:        ${\mathcal{M}}\leftarrow$ semanticProjection($\tau_{t},R^{3D}$)
13:     end if
14:     $c_{t+1}\leftarrow$ getNextBestPose($\mathcal{F},\mathcal{V},g_{k}$)
15:     goTo($c_{t+1}$)
@@
V Experimental Results
The goals of our experiments are 1) to investigate the signage recognition accuracy of our proposed scene understanding method (Sec. V-C), and 2) to evaluate our exploration-exploitation planner in terms of the average signage coverage rates and exploration time per signage (Sec. V-D).
V-A Experimental Setup
We evaluate the proposed system in two large-scale shopping malls with four and nine shops or restaurants, respectively. We conduct the experiments on the Scout-mini mobile robot platform shown in Fig. 5. The sensor suite comprises a Realsense D455 RGB-D camera and an OS0-128 LiDAR. The computing is performed on an Intel NUC mini PC with Intel Core i9-12900H and an NVIDIA Jetson AGX Orin module with 16GB GPU memory. The WIFI router is used to build a local network for communication among the robot platform, onboard computers and sensors.
@@ V-B Implementation Details
We elaborate the venue maps of the two scenarios via Google My Maps [^2] https://www.google.com/maps/d/ according to the original Google Maps to showcase the landmarks of interest. We adopt FAST-LIO2 [38] for robot pose estimation, a dynamic object removal module [7] for point cloud segmentation, and OctoMap [39] for constructing the 3D volumetric maps simultaneously. We then squeeze the 3D maps into 2D occupancy grid maps. We detect and filter the frontiers by [9], while we generate the viewpoints by heading towards the signage with a certain distance. The hyperparameters used in our method are listed in Table I.
Transformation Estimation: We adopt MAGSAC [35] together with Levenberg-Marquardt refinement to estimate the transformation matrix between the online and venue map by minimizing the sum of the squared error $E(R,t,\alpha)=\frac{1}{N}\sum_{i=0}^{N}||p_{o}-\alpha(Rp_{g}+t)||^{2}$, where $p_{o}$ denotes the centroid of the online detected landmark and $p_{g}$ refers to that on the venue map. We instantiate the above algorithms using the 2D affine transformation method  [^3] https://docs.opencv.org/3.4.1/d9/d0c/group__calib3d.html.
Figure 5: The details of the wheeled vehicle used in real-world experiments.
TABLE I: Hyperparameters
Parameter Value
$h$ num of hop neighbouring to previous landmark 1
@@
$\beta$ factor for balancing exploration and exploitation 9
@@
Dynamic Object Removal: To alleviate the influence of pedestrians during robot exploration, a dynamic object removal approach inherited from [7] is deployed, where the front-end sub-module uses a range-image-based method to coarsely detect the points from dynamic objects, followed by a static mapping framework in the back-end sub-module. Each frame from the LiDAR data stream is effectively separated into dynamic (pedestrians) and static (building structures) scan points, while only the static part is sent into OctoMap [39].
Prompts: We prompt AnyText [33] model to generate the text images by: a logo of a store, with ‘xx’ written on it, which means: {CJK}UTF8gbsn 一个店铺的logo,写着“xx” in Chinese, where xx is the landmark name extracted from venue maps using CnOCR [^4] https://github.com/breezedeus/cnocr.
V-C Signage Recognition Accuracy
@@
To test our signage understanding approach, we evaluate the signage recognition recall rates on a signage images dataset collected in the two scenarios, comparing it with four alternative methods: 1) using CLIP [26] to perform text-to-image similarity matching between landmark names and detected images, 2) using Chinese-CLIP [40] that is fine-tuned on Chinese dataset, 3) using recognition results of an OCR model, ESTextSpotter [27], with Levenshtein distance as the text-level measurement, 4) Retrieval using font: using the latent features extracted from ESTextSpotter and performing image-level retrieval with font-rendered text images [31]. Our method (Retrieval using diffusion) utilizes a text-diffusion model for text-to-image conversion compared to using font. We collect the signage images of 18 shops and restaurants with 5 different views in real-world shopping malls. The examples of signage recognition in open-world scenarios shown in Fig. 6 demonstrate the difficulty. We consider signage to be correctly recognized if the top-$k$ ($k=1,2$) retrieval result equals the ground truth label and the corresponding similarity is higher than a threshold $\gamma_{recall}=0.3$, which is decided to filter the irrelevant text that is neither written on signage nor within the venue maps during exploration.
The results are reported in Table II. It turns out that the signage recognition capabilities of both CLIP and Chinese-CLIP (ViT-B/16) are poor with only 28.9% of recall@1. We attribute it to the observation that CLIPs cannot extract effective text features from signage images that they produce very close similarities when matched with different landmark names and thus cannot filter out the irrelevant texts appearing in the scene. Moreover, the state-of-the-art STS method [27] still struggles with signage recognition, achieving only 52.2% recall rates due to the noisy recognition results. We also show that the most common text measurement, Levenshtein distance, is infeasible for square characters since it does not consider the similarity of strokes inside the characters. By converting texts to the images using font, e.g., Arial, and calculating cosine similarity with signage images [31], the recall rates are improved by 10%. Furthermore, our diffusion-based approach yields an additional 15.6% improvement over font-rending method. We think it is because the generative images are more realistic and have more contextual features, proving that our approach improves the signage recognition capability of a closed-set detector without fine-tuning. Notably, our method’s recall@2 is 10% higher than recall@1, demonstrating its better recognition capability when the search scope is narrowed to neighbouring landmarks. By adopting a 2D-to-3D fusion strategy, our method achieves more robust results with higher similarity scores. Our method can also adapt to recognizing signage in different languages by only fine-tuning the text detectors, since AnyText [33] itself supports rending various languages on the generative images.
@@
Figure 6: Examples of signage recognition. The first column is the ground truth of the texts. In the remaining columns, red boxes highlight the texts of interest, and the words in white boxes are the corresponding recognition results from ESTextSpotter [27]. The noisy results of directly employing a closed-set text spotter along with the interference from numerous text instances that are not presented on the signage make it difficult for accurate signage recognition.
V-D Signage Coverage Efficiency
We compare our signage-aware exploration algorithm with multiple rapidly exploring randomized trees (RRTs)-based exploration [9]. For a fair comparison, we also equip RRTs-based exploration with the proposed signage understanding module. We choose two starting points in two scenarios and conduct three trials at each point for two methods, respectively. We evaluate the signage coverage efficiency by the signage coverage rates and exploration time per signage with the corresponding standard variance, which are reported in Table III. The former refers to the average number of recognized signage of all the signage in the scene, and the latter is calculated as: $\frac{1}{n}\sum T(i)/S(i)$, where $n$ is the number of trials, $T(i)$ and $S(i)$ are the total exploration time until no detected frontiers and signage coverage number of the $i$th trial, respectively. It turns out that the RRTs-based exploration method detects around 1 and 1.67 of 4 signage at two starting points in scenario 1, respectively. For the larger scenario, it can only find 3 and 4 of 9 signage. In contrast, our method strikes a balance between exploration and exploitation, which contributes to improvement in coverage rates that we can cover 3 and 3.67 of 4 signage in scenario 1 and 7 and 6.67 of 9 signage in scenario 2. Furthermore, despite the exploitation costs more time, our method can correlate the recognized signage with venue maps to know the approximate direction towards the next landmark, resulting in reducing more than 1 time the exploration time per signage than the baseline. In consequence, it produces more efficient exploration paths to cover more signage during the progress (see Fig. 7).
TABLE II: Recall rates of signage recognition ($\uparrow$)
Recall@1 Recall@2
@@ V-D Signage Coverage Efficiency
Retrieval using font [31] 62.2% 73.3%
Retrieval using diffusion (Ours) 77.8% 87.8%
TABLE III: Signage coverage rates ($\uparrow$) & exploration time per signage ($\downarrow$).
@@
(4 landmarks) Coverage rates Average time ($s$) Coverage rates Average time ($s$)
@@
(9 landmarks) Coverage rates Average time ($s$) Coverage rates Average time ($s$)
@@
Some qualitative examples of the exploration trajectories are illustrated in Fig. 8. We see that RRTs-based exploration in green blindly explores the unknown boundaries, which can only successfully recognize the signage whenever the robot happens to face to the signage with proper orientations. Moreover, it often stops early before fully exploiting the regions and misses some shops at the boundaries of scenarios. In contrast, the trajectories of our method in red perform exploitation behaviour to actively spot the signage in the scene whenever its utility is higher than that of exploration. Once two signage instances are recognized, our method calculates the directional heuristic to induce the robot to reach the next landmarks rapidly and reduce redundant exploration. Besides, our method plans a landmark route on the venue maps by a TSP solver that enables the robot to traverse the signage in the scenarios with less overlapping areas. Videos of real-world experiments are given at https://youtu.be/sFSb8spZ_bo.
Figure 7: The exploration progress of two methods in four scenario settings. The curves stop at the time when the last new signage is covered, after which the exploration may still continue but no more signage is covered.
V-E Ablation study
An important design choice is the weighting parameter for balancing exploration and exploitation. Therefore, we study the influence of the balancing weight by comparing our method with different $\beta=\{3,9,15\}$ in scenario 2 starting from point 1. The higher the weight, the more importance is given to the exploitation. The experimental results are reported in Table IV and the trajectory comparison is illustrated in Fig. 9. We see that when the weight is too low ($\beta=3$), the planner inclines exploring to the frontiers under the guidance of venue maps, but delays or even misses recognizing signage, resulting in lower coverage rates. Although it can finally find out part of missed signage using the directional heuristic, it takes more time and decline the efficiency (see Fig. 9a). In contrast, as $\beta=15$, the planner biases selecting too many viewpoint candidates that it spends more time on perceiving the potential signage nearby the robot and decreases the efficiency to explore further signage (see Fig. 9b). A moderate balance of exploration and exploitation ($\beta=9$) can achieve both higher signage coverage rates and searching efficiency.
@@ V-E Ablation study
Balance weight Coverage rates Average time ($s$)
Scenario 2 starting point 1 $\beta=3$ 6.00 $\pm$ 1.00 / 9 143.17 $\pm$ 20.10
$\beta=9$ 7.00 $\pm$ 1.00 / 9 93.40 $\pm$ 14.23
$\beta=15$ 6.67 $\pm$ 0.58 / 9 123.99 $\pm$ 14.01
Figure 8: Qualitative examples of the exploration trajectories of our method and baselines in four scenario settings, showing that our method produces more efficient and reasonable exploration paths by using the venue maps.
Figure 9: Ablation experiments for evaluating the influence of balancing weight $\beta$. Triangles and circles show the views and poses of robots when exploitation for ablation methods ($\beta=3$ and $15$).
VI Conclusion
We have presented the first signage-aware exploration method that utilizes venue maps to incorporate text-level information for searching shops in open-world environments. By harnessing venue-specific text sets and retrieving the most similar text, our proposed diffusion-based text instance retrieval effectively detects and recognizes signage in the scene. Additionally, a 2D-to-3D semantic fusion strategy is employed to enhance the recognition accuracy. Moreover, we design an exploration-exploitation planner to guide exploration using venue maps while balancing the exploration of new areas with the exploitation of known spaces to recognize signage actively. Experimental results demonstrate that our method surpasses state-of-the-art scene text spotting methods and traditional exploration techniques in terms of both signage recognition accuracy and coverage efficiency. Our method is applicable to navigation tasks by setting goals directly on the venue maps.
A limitation of the proposed method lies in the situations where the text is absent. A promising direction is integrating the structure-aware localization methods [3, 4] to improve exploration efficiency. Another potential work is to enable robots to check if previous signage recognition is wrong when the current signage is not the same as the expected landmark name. Introducing shopfront detection methods [41] can also help find the shops in scenarios with irregular layouts.

### v2

@@ IV-C Exploration-Exploitation Planning
Candidate Next Pose: During navigating to the next landmark $g_{t}$, the candidate next poses $c\in V$ are sampled among the set of unvisited frontiers $\mathcal{F}$ and unvisited viewpoints $\mathcal{V}$. The frontiers $f\in\mathcal{F}$ bias the exploration to unknown regions, while the viewpoints $v\in\mathcal{V}$ allow the robot to approach and face the signage for better recognition.
Next Best Pose Selection: We tailor the frontier utility $U_{f}$ and viewpoint utility $U_{v}$ for the decision on the candidate selection, respectively. In the initial stage, we induce the robot to explore its surroundings by selecting the frontiers that maximize the information gain $\sum G$ within the camera’s FOV. After at least two landmarks are found, we align the online map with the venue map and estimate all the landmark poses $p_{g_{t}}$ at the world coordinate using random sample consensus (RANSAC) [36] and iterative closest points (ICP). Therefore, we perform the frontier-based informed search using the relative directions between landmarks on the venue maps. To this end, a directional heuristic is calculated to favor selecting the frontiers closest to the direction towards the subgoal $g_{t}$ (see Fig. 4), which is defined as the normalized inner product between the vectors from the current position $p_{t}$ to the frontier $f_{i}$ and the vector from the current position $p_{t}$ to the position of next subgoal $p_{g_{t}}$. Specifically, we first calculate the transformation matrix $T=[R|t]$ and scale $\alpha$, which are updated whenever a new landmark is found. Then, we back-project the landmark positions from the venue map to the world coordinates. As such, the directional heuristic $h(f_{i})$ is calculated as $h(f_{i})=\frac{(p_{t}-f_{i})\cdot(p_{t}-T^{-1}p_{g_{t}})}{\|p_{t}-f_{i}\|\|p_{t}-T^{-1}p_{g_{t}}\|}$, and thereby the frontier utility $U_{f}(f_{i},p_{t})$ is computed as:
@@
Figure 4: Position estimation of the next landmark by calculating the coordinate transformation between the online map and the venue map for guiding the frontier selection.
@@
where the viewpoint score $s_{v}(v_{j})=s_{\phi}(o_{\tau_{j}},o_{\tau_{j}}^{\prime})$, i.e., equals the recognition score of the signage text $\tau_{j}$ belonging to viewpoint $v_{j}$ when matched to the venue maps, and $d(p_{t},v_{j})$ is the navigation cost with a weight $\eta$. A factor $\beta$ that determines to what extent of scores we should highlight the viewpoint candidates than frontiers is employed to handle the exploration-exploitation dilemma. Intuitively, the priority of exploiting signage originates from the fact that it can help localize the robot on the venue maps and facilitate the decision on the next best pose. Finally, we obtain the optimal next pose $c_{t+1}$ among unvisited frontiers and viewpoints that maximizes the overall utility $U(c_{j},p_{t})$:
@@
where $c_{j}\in\mathcal{F}\cup\mathcal{V}$ and $I_{f}(c_{j})=1\ \text{if}\ c_{j}\in\mathcal{F}\ \text{else}\ 0$. Then we plan a collision-free path to $c_{t+1}$ using the information RRT* planner from Open Motion Planning Library [37] with a fixed replan rate, and follow the route by a local planner that excels in dynamic collision avoidance  [^1] https://www.cmu-exploration.com/. The overall procedure is summarized in Algorithm 1.
V Experimental Results
V-A Experimental Setup
We evaluate the proposed system in two large-scale shopping malls with 4 and 9 landmarks, respectively. We implement the system on a Scout-mini mobile robot platform shown in Fig. 2. The sensor suite comprises a Realsense D455 RGB-D camera and an OS0-128 LiDAR. The computing is performed on an Intel NUC mini PC with Intel Core i9-12900H and an NVIDIA Jetson Orin NX with an AI performance of 157 TOPS. The WIFI router builds a local network for communication among the robot platform, sensors, and onboard computers.
@@ V-A Experimental Setup
0:  Venue map $\mathbf{M}$ including landmark names $\mathcal{T}$ and a topological graph $\mathcal{G}$, prior features pool $\mathcal{D}_{e}$, text spotter $\phi$, RGBD image {$o_{t}$, $D_{t}$}, LiDAR $\mathcal{X}_{t}$, signage map $\mathcal{M}=\emptyset$, frontiers $\mathcal{F}=\emptyset$, viewpoints $\mathcal{V}=\emptyset$
1:  while $\mathcal{F}\neq\emptyset\ or\ t=0$ do
2:     $o_{t},D_{t},\mathcal{X}_{t}\leftarrow$ getObservation()
3:     ${\mathcal{M}},f_{t}\leftarrow$ mappingModule($\mathcal{X}_{t}$)
4:     $\mathcal{F}\leftarrow\mathcal{F}\cup f_{t}$
5:     $R^{2D},e^{2D},v_{t}\leftarrow\phi(o_{t})$
6:     if $R^{2D}$ is not None then
7:        $R^{3D},e^{3D}\leftarrow$ 3DFusion($R^{2D},e^{2D},D_{t},\mathcal{M}$)
8:        $\tau_{t}\leftarrow$ textInstanceRetrieval($e^{3D},\mathcal{D}_{e}$)
9:        $g_{t}\leftarrow$ topologicalPlanning($\tau_{t},\mathcal{G}$)
10:        $\mathcal{V}\leftarrow\mathcal{V}\cup v_{t}$
11:        ${\mathcal{M}}\leftarrow$ semanticProjection($\tau_{t},R^{3D}$)
12:     end if
13:     $c_{t+1}\leftarrow$ getNextBestPose($\mathcal{F},\mathcal{V},g_{t}$)
14:     goTo($c_{t+1}$)
15:     $t\leftarrow t+1$
@@ V-B Implementation Details
We obtain the venue maps of the two scenarios via Google My Maps [^2] https://www.google.com/maps/d/ to showcase the landmarks of interest. We use a TSP solver [38] to plan the landmark routes on the venue maps. We adopt FAST-LIO2 [39] for robot pose estimation, a dynamic object removal module [7] for point cloud segmentation, and OctoMap [40] for constructing a 3D volumetric map simultaneously. We then squeeze the 3D map into a 2D occupancy grid map. We detect and filter the frontiers by [8]. We generate the viewpoints by heading toward the signage at a proper distance. We adopt MAGSAC [36] with Levenberg-Marquardt refinement to estimate the transformation matrix between the venue map and world coordinates. The hyperparameters are listed in Table I, where $\eta$ is chosen based on the distance range of the frontiers in the scenes; $\lambda$ reduces the bias caused by viewpoints often being closer to the robot than frontiers; $\mu_{radius}$ and $\mu_{gain}$ defines the range of frontiers within the robot’s surroundings.
To alleviate the influence of pedestrians during robot exploration, a dynamic object removal approach inherited from [7] is deployed, where the front-end sub-module uses a range-image-based method to coarsely detect the points from dynamic objects, followed by a static mapping framework in the back-end sub-module. Each frame from the LiDAR data stream is effectively separated into dynamic (pedestrians) and static (building structures) scan points, while only the static part is sent into OctoMap [40].
We prompt AnyText [34] model to generate the text images by: a sign of a store with [TEXT] written on it, which means: {CJK}UTF8gbsn 一个店铺标识,写着“[TEXT]” in Chinese, where [TEXT] is a place name extracted from the venue map using CnOCR [^3] https://github.com/breezedeus/cnocr.
Hyperparamter Value
$h$ num of hop neighboring to previous landmark 1
@@
$\beta$ balance weight for exploration and exploitation 9
@@
TABLE I: Hyperparamters of the robotic system for real-world experiments
V-C Signage Recognition Performance
@@
This experiment is to investigate the signage recognition performance of our proposed signage understanding method. We evaluate our method based on signage recognition recall rates using a dataset of signage images collected in two scenarios, comparing it with four alternative methods: 1) using CLIP [27] to perform text-to-image similarity matching between landmark names and detected images, 2) using Chinese-CLIP [41], fine-tuned on Chinese dataset, 3) using recognition results of an OCR model, ESTextSpotter [28], with Levenshtein distance as the text-level measurement, and 4) font-based rendering retrieval (FontRR): leveraging the latent features extracted from ESTextSpotter and performing image-level retrieval with font-rendered text images [32]. Our method, termed diffusion-based rendering retrieval (DiffusionRR), utilizes a text-diffusion model for text-to-image generation. We collect the signage images of 18 shops and restaurants with 5 different views in real-world shopping malls. Some examples of signage recognition in open-world scenarios, shown in Fig. 5, illustrate the challenges inherent in this task. We consider a sign to be correctly recognized if the top-$k$ ($k=1,2$) retrieval results contain the ground truth and the similarity score is higher than a threshold $\gamma_{recall}=0.3$ for filtering out texts that are neither displayed on signage nor within the text set of venue maps.
@@
Figure 5: Examples of signage recognition. The first column shows the ground truths of the texts. In the remaining columns, red boxes highlight the texts of interest, and the words in white boxes are the corresponding recognition results from ESTextSpotter [28]. The noisy recognition results and the interference from text instances that are not presented on the signage make it difficult for accurate signage recognition. Our method can robustly match the real images with generative images even though all their recognized texts are not the same as the ground truths.
Methods Recall@1 Recall@2
@@
FontRR [32] 62.2% 73.3%
DiffusionRR (Ours) 77.8% 87.8%
TABLE II: Recall rates of signage recognition ($\uparrow$)
The results are reported in Table II. It turns out that the signage recognition capabilities of both CLIP and Chinese-CLIP (ViT-B/16) are poor, achieving only 28.9% of recall@1. We attribute this to the inability of CLIP models to extract text features from signage images effectively. As a result, they produce similar similarity scores when matching different landmark names, making it difficult to filter out irrelevant texts in the scene. Moreover, the state-of-the-art STS method [28] still struggles with signage recognition, achieving only 52.2% recall rates due to the noisy recognition results. We also show that the most common text measurement, Levenshtein distance, is infeasible for non-alphabet square-shaped characters since it does not consider the similarity of strokes inside the characters. By converting texts to images using font, e.g., Arial, and calculating cosine similarity with signage images [32], the recall rates are improved by 10%. Furthermore, our diffusion-based approach yields an additional 15.6% improvement over the font-rending method. We think it is because the generative images are more realistic and have more contextual features, proving that our approach improves the signage recognition capability of a closed-set detector without fine-tuning. Notably, our method’s recall@2 is 10% higher than recall@1, demonstrating better recognition capability when the matching scope is narrowed to neighboring landmarks. By adopting the 2D-to-3D fusion strategy, our method achieves more robust results with higher similarity scores.
V-D Signage Coverage Efficiency
This experiment is to evaluate the improvement in the efficiency of covering all the landmarks by using venue maps and exploitation method, comparing with Multiple Rapidly Exploring Randomized Trees (RRTs)-based exploration [8], which selects the frontiers by a utility similar to $U_{f}(f_{i},p_{t})$ but with substitution of $h(f_{i})$ by information gain $\sum G$. We also equip RRTs-based exploration with the proposed signage understanding module. We consider a landmark to be successfully covered if its corresponding signage is accurately recognized, thus here we use signage to represent landmarks. We choose two starting points in each scenario and conduct three trials at each point for two methods, respectively. We evaluate the signage coverage efficiency by both signage coverage rates and exploration time per sign with the standard variance. The former refers to the average number of recognized signs of all the signs in the scene, and the latter is calculated as $\frac{1}{n}\sum T(i)/S(i)$, where $n$ is the number of trials, $T(i)$ and $S(i)$ are the total exploration time until no detected frontiers and signage coverage number of the $i$-th trial, respectively. We first compare our entire system with the baseline [8] and then study the respective impact of venue maps and exploitation in the balance.
Figure 6: Qualitative examples of the exploration trajectories in four scenario settings. Our method produces more efficient and reasonable exploration paths by using venue maps.
@@ V-D Signage Coverage Efficiency
(4 landmarks) Coverage rates $\uparrow$ Average time ($s$) $\downarrow$ Coverage rates $\uparrow$ Average time ($s$) $\downarrow$
@@
(9 landmarks) Coverage rates $\uparrow$ Average time ($s$) $\downarrow$ Coverage rates $\uparrow$ Average time ($s$) $\downarrow$
@@
TABLE III: Coverage rates and exploration time per sign.
[…]
Figure 7: The exploration progresses in four scenarios. The curves stop when the last new sign is covered, after which the exploration may continue but no more sign is covered.
An important design choice is the weighting parameter $\beta$, which balances venue map-guided exploration and exploitation. To investigate the impact of venue maps and exploitation, we evaluate $\beta$ with values $\{3,9,15\}$. A higher weight assigns greater importance to exploitation. The experimental results tested in scenario 2 starting from point 1 are reported in Table IV and the corresponding trajectories are compared in Fig. 8. When the weight is too low ($\beta=3$), the planner is approximately degraded to the RRT-based exploration with only venue maps. While inclining to frontier exploration, it often fails to recognize certain signs due to poor observations. Although it can eventually revisit part of the missed signs using venue maps, it often requires more time for the robot to localize itself in the venue maps, which significantly decreases the utility of venue maps and the overall path efficiency (see Fig. 8a). On the other hand, as the weight is too high ($\beta=15$), the planner is approximately degraded to the RRT-based exploration with only exploitation that it prioritizes selecting viewpoints for perceiving the potential signage. While this enables the robot to cover more signage and frequently localizes itself in the venue maps, excessive exploitation constrains the robot from exploring beyond the surroundings, hindering discovering the distant signs (see Fig. 8b). A moderate balance of exploration and exploitation ($\beta=9$ by default) can achieve both higher signage coverage rates and search efficiency. Videos of real-world experiments are shown on our project website.
Figure 8: Qualitative examples for evaluating the impact of balance weight $\beta$. Triangles and circles show the views and poses when performing exploitation with $\beta=3$ and $15$.
Methods Coverage rates $\uparrow$ Average time ($s$) $\downarrow$
Scenario 2 starting point 1 [8] 3.33 $\pm$ 0.58 / 9 186.67 $\pm$ 67.79
Ours ($\beta=3$) 6.00 $\pm$ 1.00 / 9 143.17 $\pm$ 20.10
Ours ($\beta=15$) 6.67 $\pm$ 0.58 / 9 123.99 $\pm$ 14.01
Ours ($\beta=9$) 7.00 $\pm$ 1.00 / 9 93.40 $\pm$ 14.23
@@
VI Discussion
1) Computational overhead. Our system relies on lightweight algorithms optimized for real-time applications. The text detection and retrieval processes run asynchronously, consuming approximately 5GB of GPU memory and achieving a speed of $\sim$2fps on our edge device (157 TOPS). This ensures high performance in signage understanding with minimal delay. 2) Scalability. Our approach may be less effective in large scenes with sparse signage, as the lack of signs can hinder localization. While our approach emphasizes the effectiveness of leveraging textual information in the scene, one can always integrate our approach into conventional ones to leverage the spatial structures for global localization. In scenes with dense signage, while false positives may increase, our system includes a RANSAC step to eliminate spatially distant matches. Additionally, dense signage can benefit text-absent issues, aiding in initial localization and exploration. 3) Multilingualism. In our scenes, some signs display texts in multiple languages, e.g., Japanese and Korean, which can interfere with accurately recognizing target texts in Chinese and English. Our method addresses this through appearance-based similarity matching, even though our text detector has not been trained on these additional languages. Fine-tuning the text detectors on multilingual datasets can further enhance multilingual recognition, as AnyText [34] supports rendering multilingual text on generative images. 4) Limitations. Our system may fail to detect signs due to camera over-exposure. This is particularly relevant for signs made of LED lights, where careful adjustment of the exposure rate is crucial. Incorporating shopfront detection methods [42] can also help locate the shops when the signage is significantly obscured.
VII Conclusion
We present the first signage-aware exploration method that leverages signage in the scenes and the 2D non-metric venue maps to incorporate text-level information for searching for landmarks in unknown open-world environments. To overcome the challenge of signage recognition, we proposed a diffusion-based text instance retrieval method that detects and recognizes signage with arbitrary shapes and styles in the scene effectively. A 2D-to-3D semantic fusion strategy is employed to enhance recognition performance. Furthermore, we design a venue map-guided exploration-exploitation planner to achieve both high signage coverage rates and search efficiency. Real-world experiments demonstrate our method is more efficient and robust compared to the baselines.